# UD-English-CHILDES: A Collected Resource of Gold and Silver Universal Dependencies Trees for Child Language Interactions

### v1

@@ Abstract
CHILDES is a widely used resource of transcribed child and child-directed speech. This paper introduces UD-English-CHILDES, the first officially released Universal Dependencies (UD) treebank derived from previously dependency-annotated CHILDES data with consistent and unified annotation guidelines. Our corpus harmonizes annotations from 11 children and their caregivers, totaling over 48k sentences. We validate existing gold-standard annotations under the UD v2 framework and provide an additional 1M silver-standard sentences, offering a consistent resource for computational and linguistic research.
@@
Xiulin Yang1  Zhuoxuan Ju1  Lanni Bu1  Zoey Liu2  Nathan Schneider1 1Georgetown University 2University of Florida {xy236, zj153, lb1437, nathan.schneider}@georgetown.edu liu.ying@ufl.edu
@@ 1 Introduction
As UD treebanks have become valuable resources in both NLP (e.g., Jumelet et al., 2025; Opitz et al., 2025) and language acquisition research (e.g., Clark et al., 2023; Hahn et al., 2020), there have been increasing efforts to parse CHILDES data using tools such as stanza (Liu and MacWhinney, 2024). However, the resulting annotation quality is often inconsistent and cannot be guaranteed. In this paper, we compile, harmonize, and manually correct major UD-style annotations of CHILDES data into a consistent, unified UD format, resulting in a gold-standard treebank of 48K sentences and 236K tokens (including, e.g., the tree in Figure 1). In addition, we construct a larger silver-standard treebank of 1M sentences and 6M tokens and report parser accuracy estimates. We publicly release both datasets.111pseudo-link-for-childes-UD
@@
Corpus Children Gold Annotation Speakers UPOS Feats Utterances Tokens
@@
S+24 Adam Trees, UPOS; features from original CHILDES Adults Gold Convrtd 17,233 0 91,114 0
LP21 Eve Trees; others unspecified All Silver Silver 2,207 108,044 8,497 532,319
LP23 10 Children Trees; others unspecified All Silver Silver 34,530 1,101,061 168,284 6,461,084
UD-English-CHILDES 11 Children Trees, UPOS All Gold N/A 48,183 1,197,471 236,941 6,892,314
Table 1: Overview of CHILDES-based UD treebanks compiled in this paper. Source corpus labels (S+24, LP21, LP23) are defined in section 3. Note that there is overlap in the Adam data: S+24 figures are counts from the original dataset; for our version, these were filtered to avoid duplicates and merged with corresponding LP23 utterances.
@@ 2.1 CHILDES Corpora
CHILDES has played a central role in both language acquisition research and the development of NLP tools. In addition to specialized corpora—such as clinical datasets (Gillam and Pearson, 2004), naturalistic family interactions (Gleason, 1980), and controlled laboratory studies (Newman et al., 2016)—CHILDES supports a wide range of approaches to developmental linguistics. Many of its corpora inform foundational theories of language acquisition, particularly the poverty of the stimulus hypothesis (Chomsky, 1976). Researchers frequently use child-directed speech from CHILDES to quantify the distribution of linguistic structures that are central to these theories, such as wanna contraction (Getz, 2019), anaphoric one (Foraker et al., 2009; Pearl and Mis, 2011), auxiliary fronting (Perfors et al., 2011), and syntactic islands (Pearl and Mis, 2011). It has also been used in computational models of language acquisition (e.g., Abend et al., 2017).
@@
# sent_id = 17235906
# childes_toks = who's that
# corpus_name = Providence
@@
# speaker_age = 37.49358303045237
# speaker_gender = male
# speaker_role = Mother
# type = question
@@ 1 Who who PRON WP _ 0 root 0:root _
Figure 2: Example of a gold-annotated CoNLL-U sentence from the CHILDES-Providence corpus.
2.2 CHILDES Dependency Treebanks
@@
More recently, UD-style annotations have been introduced to CHILDES. Liu and Prud’hommeaux (2021) used a semi-automatic method to convert previous GR-based annotations into UD trees, focusing on child-produced speech (ages 18–27 months) from the Eve data within the Brown corpus (Brown, 1973). Subsequently, Szubert et al. (2024) developed gold-standard UD annotations by automatically transforming GR annotations and manually correcting them. Their dataset includes child-directed speech from the Adam data of the Brown corpus and the Hebrew Hagar corpus (Berman, 1990), addressing spoken-language-specific phenomena such as repetitions and non-standard vocabulary, as well as a mapping to semantics.
@@ 3.1 Annotation Pipeline
Following Liu and Prud’hommeaux (2023), we collect CHILDES corpora using the childesr [^2] https://langcog.github.io/childes-db-website/ R package and identify whether each sentence has been previously annotated with a UD tree. Sentences with existing annotations are automatically tagged with Universal Part-Of-Speech tags (UPOS) using stanza333stanza 1.9.2 (Qi et al., 2020), while unannotated sentences are assigned both UPOS and dependency trees. Our current work focuses on correcting previously human-annotated data. To ensure conformity with UD guidelines, we run all processed sentences through the UD validation tool [^4] https://github.com/UniversalDependencies/tools/blob/master/validate.py and manually fix those that fail validation. The correction work is performed by three linguistics graduate students trained in UD annotation.
@@ 3.2 Harmonization
As each treebank follows its own annotation guidelines, we performed a series of normalization steps to harmonize them into a consistent format. Our unified format is primarily based on LP23, with several adaptations described below.
@@
In our normalized CoNLL-U files, we include the following metadata fields with an example provided in Figure 2: sent_id (utterance ID retrieved via the childesr R package); childes_toks (tokenized utterance); corpus_name (original corpus name); gold_annotation (indicates whether the sentence is manually annotated); speaker_gender, speaker_role, and speaker_age (speaker metadata); text (the text aligned with the tree), and type (sentence type). Table 3 summarizes the distribution of the main sentence types and compares them with those in the UD 2.15 release of GUM (Zeldes, 2017), a multi-genre English corpus. Notably, questions occur in the CHILDES conversations at a much higher rate—they are nearly half (45%) as frequent as declarative utterances, as opposed to 9% in GUM.
Sentence Type CHILDES GUM (s_type)
declarative 31,996 7,695 (decl)
question 14,241 716 (q, wh)
imperative_emphatic 797 1,326 (imp, intj)
others 1,149 2,409
Table 3: Sentence type counts in gold CHILDES and GUM corpora. “Others” encompasses less frequent categories: trail off, interruption, self interruption, quotation next line, self interruption question, trail off question, and interruption question.
@@
Each of the three treebanks defines its own subtypes for the reparandum and parataxis relations. For example, S+24 includes parataxis:repeat, while LP23 uses parataxis:discourse, labels not present in the current UD guidelines. Similarly, LP21 and LP23 annotate reparandum with subtypes such as restart and repetition to mark special utterance features of children’s speech. To ensure consistency across treebanks, we move all such subrelation information to the MISC column.
@@
Since S+24 and LP23 overlap in the Adam corpus, we merged the annotations from these two treebanks. 3375 sentences are repetitive in S+24. We removed these sentences from our corpus777883 sentences from S+24 could not be merged and were therefore removed from our treebank as well..
To ensure a more linguistically plausible analysis, we also diverged from Liu and Prud’hommeaux (2023) in our treatment of interjections. Instead of annotating utterances consisting solely of interjections (e.g., Ha ha ha ha) as conj, we used the flat relation, as illustrated in Figure 3.
{dependency}
[edge unit distance=2ex] {deptext}[column sep=.5em] Ha & ha & ha & ha & .
[…]
{dependency}
[edge unit distance=2ex] {deptext}[column sep=.5em] Ha & ha & ha & ha & .
[…]
Figure 3: A tree example before (left) and after (right) standardization.
@@ 3.3 Silver Data Assessment
To create silver-standard annotations, we apply stanza to the unannotated portion of the corpus. To estimate the quality of these silver annotations, we evaluate the parser’s performance on the gold-standard data. We report Labeled and Unlabeled Attachment Scores (LAS/UAS) in Table 4. The parser achieves an overall LAS of 83.3. Performance is higher on parents’ speech (86.3 LAS) than on children’s speech (81.2 LAS), likely due to the greater syntactic regularity and lower frequency of disfluencies in adult utterances.
@@ 4 Conclusion & Future Work
In this paper, we present the first harmonized UD treebanks for CHILDES, covering 11 corpora and over 48k sentences from both child-directed and child-produced speech. Future work will involve further corrections to the silver data and the expansion of the treebanks. We invite collaboration on this ongoing effort.

### v3

@@ Abstract
CHILDES is a widely used resource of transcribed child and child-directed speech. This paper introduces UD-English-CHILDES, the first officially released Universal Dependencies (UD) treebank. It is derived from previously dependency-annotated CHILDES data, which we harmonize to follow unified annotation principles. The gold-standard trees encompass utterances sampled from 11 children and their caregivers, totaling over 48K sentences (236K tokens). We validate these gold-standard annotations under the UD v2 framework and provide an additional 1M silver-standard sentences, offering a consistent resource for computational and linguistic research.
@@
[…]
@@ 1 Introduction
As UD treebanks have become valuable resources in both NLP (e.g., Jumelet et al., 2025; Opitz et al., 2025) and language acquisition research (e.g., Clark et al., 2023; Hahn et al., 2020), there have been increasing efforts to parse CHILDES data using tools such as stanza (Liu and MacWhinney, 2024). However, the resulting annotation quality is often inconsistent and cannot be guaranteed. In this paper, we compile, harmonize, and manually correct major UD-style annotations of CHILDES data into a consistent, unified UD format, resulting in a gold-standard treebank of 48K sentences and 236K tokens (including, e.g., the tree in Figure 1). In addition, we construct a larger silver-standard treebank of 1M sentences and 6M tokens produced by stanza 111stanza 1.9.2 (combined model) and report parser accuracy estimates. We publicly release both datasets. [^2] Official gold UD release: https://github.com/UniversalDependencies/UD_English-CHILDES Note: Due to a postprocessing error, the gold UD release from the main branch is missing approximately 10K sentences. For complete access to the data, please use the dev branch. The main branch will be updated in the next official release scheduled for November 2025.
Silver release: https://github.com/xiulinyang/UD-CHILDES.
@@
Corpus Children Speakers Trees UPOS Feats Utterances Tokens
@@
S+24 Adam Adults Gold Gold Convrtd 17,233 0 91,114 0
LP21 Eve All Gold Silver Silver 2,207 0 8,497 0
LP23 10 Children All Gold Silver Silver 34,530 0 168,284 0
UD-English-CHILDES 11 Children All Gold/Silver Gold/Silver N/A 48,183 1,197,471 236,941 6,892,314
Table 1: Overview of CHILDES-based UD treebanks compiled in this paper and our newly-released UD-English-CHILDES treebank. Source corpus labels (S+24, LP21, LP23) are defined in section 3. Note that there is overlap in the Adam data: S+24 figures are counts from the original dataset; for our version, these were filtered to avoid duplicates and merged with corresponding LP23 utterances. The heading Gold refers to the subset of utterances for which trees and UPOS have been manually corrected according to the UD v2 framework; Silver refers to the subset with fully automatic annotations from stanza.
@@ 2.1 CHILDES Corpora
CHILDES is a collection of child–adult conversations recorded in naturalistic or laboratory settings. It has played a central role in both language acquisition research and the development of NLP tools. In addition to specialized corpora—such as clinical datasets (Gillam and Pearson, 2004), naturalistic family interactions (Gleason, 1980), and controlled laboratory studies (Newman et al., 2016)—CHILDES supports a wide range of approaches to developmental linguistics. Many of its corpora inform foundational theories of language acquisition, particularly the poverty of the stimulus hypothesis (Chomsky, 1976). Researchers frequently use child-directed speech from CHILDES to quantify the distribution of linguistic structures that are central to these theories, such as wanna contraction (Getz, 2019), anaphoric one (Foraker et al., 2009; Pearl and Mis, 2011), auxiliary fronting (Perfors et al., 2011), and syntactic islands (Pearl and Mis, 2011). It has also been used in computational models of language acquisition (e.g., Abend et al., 2017).
@@
# sent_id = 22497 (normalized sentence ID across corpora; used to avoid
collisions since some corpora share identical sentence IDs)
# original_sent_id = 946255 (original sentence ID from the corpus, as assigned
in childsr)
# childes_toks = who's that (original token string from childsr)
# child_name = Abe
# corpus_name = Kuczaj
@@
# speaker_age = 43.72369042485472 (child's age in months)
# speaker_gender = male (child's gender)
# speaker_role = Father (speaker role in conversation)
# type = question (sentence type annotation)
@@ 1 Who who PRON WP _ 0 root 0:root _
Figure 2: Example of a gold-annotated CoNLL-U sentence from the CHILDES-Providence corpus, with added parenthetical explanations of sentence-level metadata. Enhanced UD (EUD) relations are added deterministically by the script at https://github.com/amir-zeldes/gum/blob/master/_build/utils/eng_enhance.ini.
2.2 Spoken Language Treebanks
Overview
The development of UD project has fostered the development of spoken language annotations across a wide variety of languages, such as Beja (Kahane et al., 2021) and Japanese (Omura et al., 2023), as documented in Dobrovoljc (2022). For English, the GUM corpus (Zeldes, 2017) incorporates several spoken genres.
CHILDES Dependency Treebanks
@@ 2.2 Spoken Language Treebanks
More recently, UD-style annotations have been introduced to CHILDES. Liu and MacWhinney (2024) release an automatically parsed version of the English CHILDES corpus, annotated with UD trees using stanza. Liu and Prud’hommeaux (2021) used a semi-automatic method to convert previous GR-based annotations into UD trees, focusing on child-produced speech (ages 18–27 months) from the Eve data within the Brown corpus (Brown, 1973). Subsequently, Szubert et al. (2024) developed gold-standard UD annotations by automatically transforming GR annotations and manually correcting them. Their dataset includes child-directed speech from the Adam data of the Brown corpus and the Hebrew Hagar corpus (Berman, 1990), addressing spoken-language-specific phenomena such as repetitions and non-standard vocabulary, as well as a mapping to semantics.
@@
Although Liu and Prud’hommeaux (2021, 2023) provide manually corrected UD trees, their annotations are inconsistent with the UD v2 framework, lack Universal Part-of-Speech (UPOS) tags, and have not been independently verified. Szubert et al. (2024) offer verified data, but they follow the UD v1 annotation guidelines. To date, there is no official UD release for CHILDES speech data.
@@ 3.1 Data Source & Statistics
In the official UD release, we divide the corpus based on the children’s names and genders. The training and dev splits (90% and 10%, respectively) are constructed from the data of Adam, Lily, Naima, Sarah, Roman, Laura, and Abe. The corpora of Eve, Violet, Emma, and Thomas are reserved for the test split. Details are reported in Table 3.
Split Children Corpus Gold Sents
Train Adam, Lily, Naima, Sarah, Roman, Laura, Abe Brown, Providence, Weist, Kuczaj, Braunwald 34,732
Dev Adam, Lily, Naima, Sarah, Roman, Laura, Abe Brown, Providence, Weist, Kuczaj, Braunwald 3,860
Test Eve, Violet, Emma, Thomas Brown, Providence, Weist, Thomas 9,591
Table 3: Data splits for the official UD_English-CHILDES with associated children, corpora, and gold-standard sentence counts.
@@ 3.2 Annotation Pipeline
Following Liu and Prud’hommeaux (2023), we collect CHILDES corpora using the R package childesr (Sanchez et al., 2019). [^3] https://langcog.github.io/childes-db-website/ Sentence normalization can be found in the paper. As the data from LP21 and LP23 are only parsed but not tagged yet, sentences with existing dependency annotations are identified and automatically tagged with UPOS using stanza (Qi et al., 2020), while unannotated sentences are assigned both UPOS and dependency trees. Our current work focuses on correcting previously human-annotated data. To ensure conformity with UD guidelines, we run all processed sentences through the UD validation tool [^4] https://github.com/UniversalDependencies/tools/blob/master/validate.py and manually fix those that fail validation. The correction work is performed by three linguistics graduate students trained in UD annotation. In total, we made approximately 8,000 corrections.
@@ 3.3 Harmonization
Each treebank follows its own annotation guidelines, which are largely based on UD but not fully compliant. We performed a series of normalization steps to harmonize them into a consistent format. Our unified format is primarily based on LP23, with several adaptations described below.
@@
In our normalized CoNLL-U files, we include the following metadata fields with an example provided in Figure 2: sent_id (normalized sentence IDs); original_sent_id (utterance ID retrieved via the childesr R package); childes_toks (tokenized utterance); corpus_name (original corpus name); gold_annotation (indicates whether the sentence is manually annotated); speaker_gender, speaker_role, and speaker_age (speaker/child metadata); text (the text aligned with the tree), and type (sentence type). Table 4 summarizes the distribution of the main sentence types and compares them with those in the UD 2.15 release of GUM (Zeldes, 2017), a multi-genre English corpus. Notably, questions occur in the CHILDES conversations at a much higher rate—they are nearly half (45%) as frequent as declarative utterances, as opposed to 9% in GUM.
Type CHILDES GUM
CDS CS Overall Overall
declarative 16,112 15,884 31,996 7,695 (decl)
question 2,882 11,413 14,295 716 (q, wh)
imperative emphatic 509 288 797 1,326 (imp, intj)
others 601 494 1095 2,409
Table 4: Sentence type counts in gold CHILDES and GUM corpora. Question includes question, self interruption question, trail off question, and interruption question. Others encompasses less frequent categories: trail off, interruption, self interruption, and quotation next line.
@@
Each of the three treebanks defines its own subtypes for the reparandum and parataxis relations. For example, S+24 includes labels such as parataxis:repeat not present in the current UD guidelines. Similarly, LP21 and LP23 annotate reparandum with subtypes such as restart and repetition to mark special utterance features of children’s speech. To ensure consistency across treebanks, we move all such subrelation information to the MISC column.
@@
Since S+24 and LP23 overlap in the Adam corpus, we merged the annotations from these two treebanks. 3375 sentences are repetitive in S+24. We removed these sentences from our corpus777883 sentences from S+24 could not be merged because S+24 and LP21 are using different data sources, and were therefore removed from our treebank as well..
To ensure a more linguistically plausible analysis, we also diverged from Liu and Prud’hommeaux (2023) in our treatment of interjections. Instead of annotating utterances consisting solely of interjections (e.g., Ha ha ha ha) as conj, we used the flat relation.
@@ 3.4 Silver Data Assessment
To create silver-standard annotations, we apply stanza to the utterances that were not sampled by the previous treebanks (but were from the same CHILDES datasets, i.e. conversations involving the 11 children in Table 2). To estimate the quality of these silver annotations, we evaluate the parser’s performance on the gold-standard data. We report Labeled and Unlabeled Attachment Scores (LAS/UAS) in Table 5. The parser achieves an overall LAS of 83.3. Performance is higher on parents’ speech (86.3 LAS) than on children’s speech (81.2 LAS), likely due to the greater syntactic regularity and lower frequency of disfluencies in adult utterances. The overall high-quality data can be more easily verified by human annotators than annotated from scratch. It also provides valuable training data for improving parsers on spoken language.
@@ 4 Conclusion & Future Work
In this paper, we present the first harmonized UD treebanks for CHILDES, covering 11 corpora and over 48k sentences from both child-directed and child-produced speech. The three datasets we compiled do not preserve conversational structure, and as a result, the finalized gold-standard treebank lacks coherent dialogue sequencing. Preserving such structure would require additional manual annotation to make sure all sentences are gold. However, since our annotations include the original_sent_id field, reconstructing the conversation structure is straightforward. Furthermore, morphological features have not been annotated or independently verified. Future work will focus on further corrections to the silver-standard data and the continued expansion of the treebanks. We welcome collaboration on this ongoing effort.
Acknowledgments
We acknowledge Ida Szubert, Omri Abend, Samuel Gibbon, Louis Mahon, Sharon Goldwater, Mark Steedman, and Emily Prud’hommeaux for their contributions to the original UD treebanking efforts. We also thank Brian MacWhinney for helpful discussions and anonymous reviewers for their suggestions.